What is the probability that you get a tail on the coin given you got the number 5 on the die?
It is 0.5 becasue the events are (or should be) independent.
What is the probability that 2 out of 15 students will have the same birthday?
P( 2 equal birthday in group of 15 ) ≈ 22.3%
The probability that 2 persons and only 2 persons in a random group of n persons can be calculated using the following expression*:
P(2 equal bd in group of n persons) = nC2(1-1/1461)(4/1461)Π2n-1[1-4(i-1)/1461] + nC2(1/1461)2Π2n-1[1-(4i-3)/1461]
with n=15, 30C2 = 105,
P(2 eq bd in group of 15) = 105(1-1/1461)(4/1461)[1-4/1461]
[1-8/1461][1-12/1461]∙∙∙[1-52/1461]
P(...) = 0.223263549... ≈ 22.3%
*You find this expression discussed in question "What is the probability that in a room of 8 people 2 have the same birthday ?". A simpler form of this expression that neglects February 29 of the leap day that gives a good
approximation is:
P(2 eq bd in group of n persons) = nC2(1/365)Π1n-1[1-(i-1)/365]
for n=15, this expression gives P(...) = 0.214918798... ≈ 21.5%
the probability of the offspring for a second generation.
What are the chances of being pregnant with twins if your husband is an identical twin himself?
The same as anyone else. No genetic marker has been found yet related to identical twins.
And how does this relate to coins?
What is the probability of drawing a number card from a standard 52 deck cards?
If we're talking cards of all suits 2-10 then there are 36 number cards. That would then be 36/52 which would be 69.2 percent of the time.
TRUE
What dose theoretical probability mean?
It means that you think about it but you don't actually do it.
What is the probability of rolling a sum of six or doubles?
It is 10/36 = 5/18 in a single throw of a pair of dice.
What is a probability is based on repeated trials of an experiment?
It is empirical (or experimental) probability.
What do you mean by Statistics?
The mathematics of the collection, organization, and interpretation of numerical data, especially the analysis of population characteristics by inference from sampling.
Read more: statistics
the odds theoretically are almost infinity to 1 as they could be any color. -------------------------------------------------------------------------------------------
2nd opinion
Let's say we have 3 boxes with 4 balls each.
Box A has 4 white balls.
Box B has 3 white balls.
Box C has 2 white balls.
The probability of drawing 2 W balls from;
Box A, P(2W│A)=
(4/4)∙(3/3)=
1
Box B, P(2W│B)=
(3/4)∙(2/3)=
1/2
Box C, P(2W│C)=
(2/4)∙(1/3)=
1/6
Say the probability of picking any of the 3 boxes is the same, we have;
P(A)=
1/3
P(B)=
1/3
P(C)=
1/3
Question is, given the event of drawing 2 W balls from a box taken blindly
from the 3 choices, what is the probability that the balls came from box A,
P(A│2W).
Recurring to Bayes Theorem:
P(A│2W)=
[P(A)P(2W│A)]/[P(A)P(2W│A)+P(B)P(2W│B)+P(C)P(2W│C)]=
[(1/3)(1)]/[(1/3)(1)+(1/3)(1/2)+(1/3)(1/6)]=
6/10=
0.60=
60%
P(A│2W)=
0.60=
60%
Does a normal probability distribution include a bimodal distribution?
No, the normal distribution is strictly unimodal.
Yes, you will probably ejaculate at night during sleep. It will be sticky and clearer than urine.
In a multiple choice test what is the probability of getting two answers correct?
That depends on how many questions there are, how many choices are listed for each question,
and whether any obviously-stupid answers are included among the choices. If any of those factors
changes, then the probability changes.
One thing we can guarantee, however, even without knowing any of these factors:
If you have studied the subject and know the material, then your probability of getting
correct answers increases dramatically.
What is the probability that a tossed coin will come up tails twice in a row?
Each one is 50%
Together it's 25% (.5 x .5 = .25)
What are the odds of getting shot at school?
Take everyone shot at school
divided by everyone that has gone to school
and you have a rough estimate of your odds of being shot at school, which is most likely extremely small.
since there is only one Ace of Spades and 52 cards in a deck the probability would be 1/52.